AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:13:33 2025
Author
*
:
Subject
*
:
> > 12.20 CARME IE514 4.1e-11mbar, ambient temperature 16.7 deg C > > FEE64 cabling complete > > WR source - VETAR2 borrowed from Nik Kurz and opwerating standalone ( no BuTIS reference ) > All MACBs mode 0x3 > > > 15.20 Power ON FEE64s #1 - #15 > > aida02 did not boot - /etc/dhcp/dhcpd.conf typo - fixed > > aida04 did not boot - AIDA FEE64 PSU output #4 fuses open circuit - fuses replaced from AIDA FEE64 PSU spare channel ( aida16 not connected yet ) - fixed > > aida07 did not boot - power OK - FEE64 comfort LEDs OK - network switch link LED OK - network config? > > aida08 boots - power OK - comfort LEDs OK - network switch link LED OK - requests RFS / but does not request RFS MIDAS - can ping aida08 OK, cannot telnet aida08 'connection refused' > > aida09 did not boot - FEE64 comfort LEDS OFF - swap FEE64 #9 & #10 power cables - aida09 boots OK, aida10 does not boot > Restore #9 & #10 power cables - swap AIDA FEE64 outputs #1 & #2 ( aida09 & aida10 ) - aida09 boots OK > Conclude AIDA FEE64 PSU ( no label ) output #1 is faulty. Replace AIDA FEE64 PSU ( label ISOL3 ). aida09 & aida10 boot OK > > aida15/aida16 MAC addresses were swapped in https://elog.ph.ed.ac.uk/CARME/467 - fixed > > aida01-aida06, aida09-aida15 boot OK - aida07 does not boot, aida08 does not complete boot > > 16.25 Configure MIDAS DEE64 servers aida01 - aida06, aida10 - aida15 > > FEE64 temperatures - attachment 1 > aida06 looks low ( c. 25 deg C cf. c. 40 deg C ) > > System wide checks - attachments 2-8 > aida02 fails WR decoder > > WR timestamps ( attachments 7 & 8 ) look odd/low - values always c. 0x2??? > > > 17.30 CARME IE514 OFF > > WR NIM emulator SMA cables to root MACB ( rear panel ) disconnected at front panel > > Power cycle FEE64s > > System wide checks & WR timestamp control - attachments 9-14 > no change > > 17.45 Detector bias -20, -100 & 150V - attachments 15-17 > CAEN N1419ET non-floating outputs ( to be checked ), HV cable SHV - 2x Lemo 00.250, LK1 not fitted > > 18.40 Detector bias -150V attachment 18 > > > > > To Do > Check ASIC settings > Configure HV > Check LKs > Ground cabling >
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown